# MovieDreamer: Hierarchical Generation for Coherent Long Visual Sequences

### v2

@@ 4.3 Comparison with the State-of-the-Art
Many existing story generation methods focus on fine-tuning with small datasets, exhibiting poor generalization. Consequently, we compare our method only with those that demonstrate high generalization capabilities, namely StoryDiffusion (Zhou et al., 2024) and StoryGen (Liu et al., 2023a). As shown in Figure 4, StoryDiffusion consumes significant computational overhead and tends to generate generic faces that do not align with the desired character. Similarly, StoryGen fails to preserve consistency and generates abnormal results. In contrast, our method achieves generating extremely long content while preserving both short-term and long-term consistency across multiple characters. The observation is also evidenced by the quantitative results in Table 1, where our method achieves high scores in both LT and ST. Moreover, the higher CLIP score reflects that our generated results align well with the storyline. Better IS, AS, and FID scores demonstrate our method achieves high-quality images. We showcase more qualitative results in Appendix E.
@@ Appendix A Implementation Details
We use the pretrained CLIP image encoder (Radford et al., 2021) to generate spatial tokens. The spatial tokens are compressed into two tokens using the compressor, a network with 6 Transformer blocks. The architecture of the compressor is shown in Figure 9. First, we process the spatial tokens using three transformer blocks. We then interchange the dimensions of the hidden states and compress them using a linear layer. The compressed hidden states are further processed through three transformers to obtain the final compressed token. We find that setting the compressed token to 2 is sufficient for well reconstructing the image.
Subsequently, we employ the pretrained U-Net of SDXL (Podell et al., 2023) as our decoder to render the compressed tokens back into images. Notably, the original SDXL architecture demands an additional projected text embedding as input. To address this, we concatenate and project the compressed token into the desired input space of SDXL simply using a linear layer.
Compared to the image dataset, our collected movie dataset may lack diversity in all kinds of objects in the world, which somewhat limits the model’s performance. To ensure the model develops a general understanding of various objects first, we employed a three-stage training strategy to train a more effective model. In the first stage, we fine-tuned the pre-trained U-Net and the compressor on the large existing image dataset at a resolution of 768 x 768 with a cosine scheduling learning rate that ranges from 1e-4 to 2e-5 and an AdamW optimizer for 120k steps, allowing the model to understand various objects. Our dataset comprises images from OpenImages (Kuznetsova et al., 2018), JourneyDB (Sun et al., 2023a), and Object365 (Shao et al., 2019).
@@
To help the model distinguish the script for different frames, we inserted two special tokens, [START_OF_FRAME] and [END_OF_FRAME], at the beginning and end of each frame’s script, respectively. Furthermore, in our setting, there are three different modalities: face embedding, compressed tokens, and description embedding. The face embedding is extracted with FARL (Zheng et al., 2022), and for each frame, we extract the top 5 characters. The compressed tokens are generated by leveraging our trained encoder $\mathcal{E}$, with the target frame as input. The description is divided into different sentences, and each sentence is encoded into one [CLS] token using LongCLIP (Zhang et al., 2024). We treat the remaining text such as ”Characters:” and ”Scene Elements:” as identifiers, which are tokenized the same as LLaMA, as shown in Figure 2.
@@
To enhance our model’s ability to maintain character consistency, we additionally input the description embedding encoded by LongCLIP (Zhang et al., 2024) and the face embedding extracted by FARL (Zheng et al., 2022) into the trained decoder. Both the text embedding and face embedding are aligned with the decoder’s input using two-layer MLPs and then concatenated with the compressed tokens to serve as the input for cross-attention.
@@
The $k$-mixutre GMM models the distribution $p(\mathbf{e}_{t}|\mathbf{H}_{<t},\mathbf{M}_{t})$, where $\mathbf{e}_{t}$ is the continuous real-valued image tokens of sequence length $S$ and dimension $d$. $\mathbf{H}_{<t}$ and $\mathbf{M}_{t}$ refer to the previous history information and the current multi-modal script. The GMM consists of $k$ multivariate Gaussian distribution components that are independent across the channel dimension $d$. To construct the GMM, we replace the logit prediction layer of our autoregressive model initialized using LLaMA with a linear layer with $2kd+k$ dimension for output. Subsequently, the mean $m$ and the variance $v$ of shape $S\times kd$ and mixture weights $w$ of shape $S\times k$ are derived. $m$ consists of a stack of tensors $[m^{(1)},m^{(2)},\dots m^{(k)}]$, where $m_{i}$ is of shape $S\times d$. Similarly, $v=[v^{(1)},v^{(2)},\dots v^{(k)}]$. The mixture weight is normalized along the channel dimension. Given the ground truth image embedding $\mathbf{e}$, we aim to minimize its negative log-likelihood within the GMM:
@@ Appendix B Evaluation Metrics
The Aesthetic Score (Schuhmann et al., 2022) is calculated based on the CLIP image features. Each image is first passed through the CLIP image encoder to obtain an embedding. This embedding is then fed into an additional small neural network, which predicts an aesthetic score.
@@ Appendix C Multi-modal Scripts Data Pipeline Details
In this chapter, we explain in detail how we construct the multimodal script data. Specifically, We collect a massive movie dataset, with movies coming from Condensed Movies (Bain et al., 2020) and others collected using the same methodology as Condensed Movies. The data consists of around 23,000 long videos, which subsequently results in 5 million keyframes with corresponding script annotations. The overall multimodal script synthesis is as follows:
@@
  • Extract keyframes from long videos at 1-second intervals, removing the beginning and end segments. These segments primarily contain sponsor messages, acknowledgments, and advertisements, lacking useful information. Subsequently, center-crop and resize each frame to a width of 896 pixels and a height of 512 pixels. Finally, leverage a Vision Language Model (Lu et al., 2024) to generate descriptions for the entire image.
  • Using GroundingDINO (Ren et al., 2024a) with ”Person” as the prompt, extract the top 5 characters from each frame. If the number of characters exceeds 5, the extra characters are ignored. If the number of characters is less than 5, there is no additional processing.
  • Employ an existing Vision Language Model (Lu et al., 2024) to obtain captions for each segmented character. The face of each character is extracted to construct the face ID bank.
@@
  • However, the character face IDs for each frame are derived from the same frame, which can lead to the model overfitting by directly copy-pasting characters into the generated results. To address this, we use FARL (Zheng et al., 2022) to randomly retrieve the same character’s ID from the face ID bank as input. This strategy effectively mitigates the overfitting issue.
@@ E.2 Additional Ablations on the Number of Compressed Tokens
Existing work Paint-by-Example (Yang et al., 2022) demonstrates that a single token can effectively reconstruct the original image. Building on this discovery, we added a few additional tokens. We compared the reconstructed images using 8 tokens and 2 tokens, with the metrics shown in the accompanying table below. The metrics indicate that increasing the number of tokens does not offer significant improvements in results. As shown in Table 3, we calculate the inception score, aesthetic score as well as FID score of images of different numbers of tokens. We further compare the CLIP similarity between reconstructed images and target images. The evaluation metrics demonstrate that the results of 8 tokens are slightly improved. But considering our objective of long content generation, this slight improvement is not enough for us to prioritize it when making trade-offs regarding sequence length and our limited computational resources, as 8 tokens will lead to a much shorter sequence length and demand more computational overhead. Consequently, we eventually chose 2 tokens. But we believe that if there are enough computational resources, 8 tokens will be a better choice.

### v3

@@ 4.3 Comparison with the State-of-the-Art
Many existing story generation methods focus on fine-tuning with small datasets, exhibiting poor generalization. Consequently, we compare our method only with those that demonstrate high generalization capabilities, namely StoryDiffusion (zhou2024storydiffusion) and StoryGen (liu2024intelligent). As shown in Figure 4, StoryDiffusion consumes significant computational overhead and tends to generate generic faces that do not align with the desired character. Similarly, StoryGen fails to preserve consistency and generates abnormal results. In contrast, our method achieves generating extremely long content while preserving both short-term and long-term consistency across multiple characters. The observation is also evidenced by the quantitative results in Table 1, where our method achieves high scores in both LT and ST. Moreover, the higher CLIP score reflects that our generated results align well with the storyline. Better IS, AS, and FID scores demonstrate our method achieves high-quality images. We showcase more qualitative results in Appendix E.
@@ Appendix A Implementation Details
We use the pretrained CLIP image encoder (radford2021learning) to generate spatial tokens. The spatial tokens are compressed into two tokens using the compressor, a network with 6 Transformer blocks. The architecture of the compressor is shown in Figure 9. First, we process the spatial tokens using three transformer blocks. We then interchange the dimensions of the hidden states and compress them using a linear layer. The compressed hidden states are further processed through three transformers to obtain the final compressed token. We find that setting the compressed token to 2 is sufficient for well reconstructing the image.
Subsequently, we employ the pretrained U-Net of SDXL (podell2023sdxl) as our decoder to render the compressed tokens back into images. Notably, the original SDXL architecture demands an additional projected text embedding as input. To address this, we concatenate and project the compressed token into the desired input space of SDXL simply using a linear layer.
Compared to the image dataset, our collected movie dataset may lack diversity in all kinds of objects in the world, which somewhat limits the model’s performance. To ensure the model develops a general understanding of various objects first, we employed a three-stage training strategy to train a more effective model. In the first stage, we fine-tuned the pre-trained U-Net and the compressor on the large existing image dataset at a resolution of 768 x 768 with a cosine scheduling learning rate that ranges from 1e-4 to 2e-5 and an AdamW optimizer for 120k steps, allowing the model to understand various objects. Our dataset comprises images from OpenImages (kuznetsova2018open), JourneyDB (sun2023journeydb), and Object365 (Shao_2019_ICCV).
@@
To help the model distinguish the script for different frames, we inserted two special tokens, [START_OF_FRAME] and [END_OF_FRAME], at the beginning and end of each frame’s script, respectively. Furthermore, in our setting, there are three different modalities: face embedding, compressed tokens, and description embedding. The face embedding is extracted with FARL (zheng2022general), and for each frame, we extract the top 5 characters. The compressed tokens are generated by leveraging our trained encoder $\mathcal{E}$, with the target frame as input. The description is divided into different sentences, and each sentence is encoded into one [CLS] token using LongCLIP (zhang2024longclip). We treat the remaining text such as ”Characters:” and ”Scene Elements:” as identifiers, which are tokenized the same as LLaMA, as shown in Figure 2.
@@
To enhance our model’s ability to maintain character consistency, we additionally input the description embedding encoded by LongCLIP (zhang2024longclip) and the face embedding extracted by FARL (zheng2022general) into the trained decoder. Both the text embedding and face embedding are aligned with the decoder’s input using two-layer MLPs and then concatenated with the compressed tokens to serve as the input for cross-attention.
@@
The $k$-mixutre GMM models the distribution $p(\mathbf{e}_{t}|\mathbf{H}_{<t},\mathbf{M}_{t})$, where $\mathbf{e}_{t}$ is the continuous real-valued image tokens of sequence length $S$ and dimension $d$. $\mathbf{H}_{<t}$ and $\mathbf{M}_{t}$ refer to the previous history information and the current multi-modal script. The GMM consists of $k$ multivariate Gaussian distribution components that are independent across the channel dimension $d$. To construct the GMM, we replace the logit prediction layer of our autoregressive model initialized using LLaMA with a linear layer with $2kd+k$ dimension for output. Subsequently, the mean $m$ and the variance $v$ of shape $S\times kd$ and mixture weights $w$ of shape $S\times k$ are derived. $m$ consists of a stack of tensors $m^{(1)},m^{(2)},\dots m^{(k)}]$, where $m_{i}$ is of shape $S\times d$. Similarly, $v=[v^{(1)},v^{(2)},\dots v^{(k)}]$. The mixture weight is normalized along the channel dimension. Given the ground truth image embedding $\mathbf{e}$, we aim to minimize its negative log-likelihood within the GMM:
@@ Appendix B Evaluation Metrics
The Aesthetic Score (schuhmann2022laion5bopenlargescaledataset) is calculated based on the CLIP image features. Each image is first passed through the CLIP image encoder to obtain an embedding. This embedding is then fed into an additional small neural network, which predicts an aesthetic score.
@@ Appendix C Multi-modal Scripts Data Pipeline Details
In this chapter, we explain in detail how we construct the multimodal script data. Specifically, We collect a massive movie dataset, with movies coming from Condensed Movies (DBLP:conf/accv/BainNBZ20) and others collected using the same methodology as Condensed Movies. The data consists of around 23,000 long videos, which subsequently results in 5 million keyframes with corresponding script annotations. The overall multimodal script synthesis is as follows:
@@
  • Extract keyframes from long videos at 1-second intervals, removing the beginning and end segments. These segments primarily contain sponsor messages, acknowledgments, and advertisements, lacking useful information. Subsequently, center-crop and resize each frame to a width of 896 pixels and a height of 512 pixels. Finally, leverage a Vision Language Model (lu2024deepseekvl) to generate descriptions for the entire image.
  • Using GroundingDINO (ren2024grounded) with ”Person” as the prompt, extract the top 5 characters from each frame. If the number of characters exceeds 5, the extra characters are ignored. If the number of characters is less than 5, there is no additional processing.
  • Employ an existing Vision Language Model (lu2024deepseekvl) to obtain captions for each segmented character. The face of each character is extracted to construct the face ID bank.
@@
  • However, the character face IDs for each frame are derived from the same frame, which can lead to the model overfitting by directly copy-pasting characters into the generated results. To address this, we use FARL (zheng2022general) to randomly retrieve the same character’s ID from the face ID bank as input. This strategy effectively mitigates the overfitting issue.
@@ E.2 Additional Ablations on the Number of Compressed Tokens
Existing work Paint-by-Example (yang2022paint) demonstrates that a single token can effectively reconstruct the original image. Building on this discovery, we added a few additional tokens. We compared the reconstructed images using 8 tokens and 2 tokens, with the metrics shown in the accompanying table below. The metrics indicate that increasing the number of tokens does not offer significant improvements in results. As shown in Table 3, we calculate the inception score, aesthetic score as well as FID score of images of different numbers of tokens. We further compare the CLIP similarity between reconstructed images and target images. The evaluation metrics demonstrate that the results of 8 tokens are slightly improved. But considering our objective of long content generation, this slight improvement is not enough for us to prioritize it when making trade-offs regarding sequence length and our limited computational resources, as 8 tokens will lead to a much shorter sequence length and demand more computational overhead. Consequently, we eventually chose 2 tokens. But we believe that if there are enough computational resources, 8 tokens will be a better choice.
@@
E.3 Additional Ablations on Continuous Token Supervision
We conduct additional experiments to further explore the importance of continuous token supervision as well as the L1 and L2 losses. The evaluation metrics are shown in Table 4. As visualized in Figure 22, when the model is trained using only continuous token supervision as the objective function, it tends to produce poor results with characters placed at the edges of the image. However, this problem is effectively resolved when L1 and L2 are introduced as additional objective functions during training. Our experiments demonstrate that both continuous token supervision and the L1 and L2 losses play crucial roles in optimizing the model.
Figure 22: Comparison between using only continuous token supervision and incorporating both L1 and L2. The results show that relying solely on continuous token supervision leads the model to place characters at the edges of the image, resulting in low-quality keyframes. With L1 and L2 only, the model is able to learn layout information but struggle to maintain target ID. With continuous token supervision, L1 and L2, the model is able to generate results with reasonable layout and preserve the target ID effectively.
Table 4: Quantitative comparisons of results trained with different losses.
[…]
Con-L1-L2 19.584 9.698 6.093 2.043 0.646 0.814
Con-L1-L2-ref 20.071 9.842 6.288 1.912 0.701 0.893
L1-L2 18.833 8.609 5.893 2.714 0.606 0.755
L1-L2-ref 19.431 8.774 5.979 2.560 0.616 0.776
Con 16.882 7.463 5.352 3.283 0.576 0.701
Con-ref 17.039 7.580 5.411 3.136 0.571 0.708
@@ Appendix F Limitations and Future Works
Acknowledgements
This work was supported by the National Key R&D Program of China (NO.2022ZD0160101) and the National Natural Science Foundation of China (No. 62206244).